A fuel mixture used in the early days of rocketry is composed of two liquids, hydrazine(N2H4) and dinitrogen tetraoxide(N2O4), which ignite on contact to form nitrogen gas and water vapor

if 264g of N2H4 and 106g of N2O4 are mixed, calculate the number of moles of the excess reactant


1
Expert's answer
2021-10-30T01:15:20-0400

2N2H4 + N2O4 → 3N2 + 4H2O

MM(N2H4) = 32.04 g/mol

n(N2H4) "= \frac{264}{32.04}=8.239 \;mol"

MM(N2O4) = 92.01 g/mol

n(N2O4) "= \frac{106}{92.01}=1.152 \;mol"

According to the reaction for each mole of N2O4 we need two moles of N2H4. So, N2O4 is a limiting reactant.

"n(N_2H_4)_{used} = 2n(N_2O_4) = 2 \times 1.152 = 2.304 \;mol"

The number of moles of the excess reactant (N2H4) = 8.239 -2.304 = 5.935 mol